# Better Gaussian Mechanism using Correlated Noise

### v1

@@ 3 Algorithm
By definition of the function $g$ we have that $g(X)_{i}=2\cdot f(X)_{i}-n$ for all $i\in[d]$ and $g(X)_{d+1}=Cn=d^{1/4}n$. Let $\tilde{x},\tilde{n}$ denote the output of $\mathcal{M}(X)$. Then for $i\in[d]$ we estimate the $i$’th value of $g(X)$ as $h(\mathcal{M}(X))_{i}=2\cdot\tilde{x}_{i}-\tilde{n}=2\cdot(f(X)_{i}+\eta+z_{i})-(n+2\eta)=g(X)+2z_{i}$. Since $z_{i}\sim\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}})$ we have that $2\cdot z_{i}$ and $Z_{i}$ are both distributed as $\mathcal{N}(0,\frac{d+\sqrt{d}}{\mu^{2}})$. We can estimate the final entry of $g(X)$ as $h(\mathcal{M}(X))_{d+1}=d^{1/4}\tilde{n}=d^{1/4}(n+2\eta)=g(X)_{d+1}+d^{1/4}2\eta$. Both $d^{1/4}2\eta$ and $Z_{d+1}$ are distributed as $\mathcal{N}(0,\frac{d+\sqrt{d}}{\mu^{2}})$. Since the random variables of all these estimates are independent $h(\mathcal{M}(X))$ is distributed as $\mathcal{M}^{\prime}(X)$.
Similarly, if $\tilde{y}$ denotes the output of $\mathcal{M}^{\prime}(X)$ we can estimate $f(X)$ and $n$ using the inverse of $h$. We estimate $n$ as $h^{-1}(\mathcal{M}^{\prime}(X))_{d+1}=\tilde{y}_{d+1}/d^{1/4}=(g(X)_{d+1}+Z_{d+1})/d^{1/4}=n+Z_{d+1}/d^{1/4}$ and estimate the $i$’th entry of $f(X)$ as $h^{-1}(\mathcal{M}^{\prime}(X))_{i}=\tilde{y}_{i}/2+\tilde{y}_{d+1}/(2d^{1/4})=(g(X)_{i}+Z_{i})/2+(g(X)_{d+1}+Z_{d+1})/(2d^{1/4})=f(X)_{i}+Z_{i}/2+Z_{d+1}/(2d^{1/4})$. As such, $h^{-1}(\mathcal{M}^{\prime}(X))$ is distributed as $\mathcal{M}(X)$ where $Z_{d+1}/(2d^{1/4})$ is the correlated noise added to all entries. ∎
@@
The mechanism is Algorithm 1. The properties follows from Lemmas 6, 8, and 9. ∎
Algorithm 1 relies on $d+1$ independent samples from Gaussian distributions. Alternatively, we can view the mechanism as outputting an estimate of $f(X)$ and $n$ by sampling a $d+1$ dimensional multivariate Gaussian distribution. The entries of the covariance matrix contain one of $4$ distinct values depending on (1) whether or not it is on the diagonal and (2) whether or not it is in the final row or column. This representation might be useful for analysts for estimating the effect of noise.
Corollary 11.
Let $X$, $n$, and $f(X)$ be defined as in Theorem 10 except that $f(X)_{d+1}=n$. Then the mechanism that outputs $f(X)+Z$ where $Z\sim\mathcal{N}(0,\Sigma/\mu^{2})$ satisfies $\mu$-GDP under the add/remove neighboring relation where the covariance matrix $\Sigma\in\mathbb{R}^{(d+1)\times(d+1)}$ has the following structure
@@
In this section we have presented multiple representations of our mechanism. In Lemma 9 we showed that Algorithm 1 is equivalent to applying the standard Gaussian mechanism using a different representation of the dataset which we denote $\hat{X}$. However, both of these mechanisms are defined for arbitrary real-valued input and output, and as such any implementation of either representation must be an approximation of the idealized algorithm.
Real values are often approximated in computers using floating-point representation and it is well-known that naive implementations can break the formal privacy guarantees [Mir12]. Additive noise mechanisms such as the Laplace mechanism [DMNS06] and the Gaussian mechanism [DKM+06] are typically defined as releasing the sum of a non-private query output and a sample from a distribution. This can be problematic if the mechanism is implemented using floating-point arithmetics because the output domains can differ between neighboring datasets (see [HDH+22, Theorem 2]). [HDH+22] present a technique for mitigating this issue for mechanisms that add a single noisy value to the true output.
However, in line 1 of Algorithm 1 we add two noise samples to the query output which likely causes similar issues if floating-point representations are used. To the best of our knowledge, it has not been studied whether implementations of mechanisms that rely on elliptical Gaussian noise suffer from floating-point vulnerabilities. As such, implementing the version of our mechanism from Lemma 8 might be preferred for production code. The values of $f(X)$ and $n$ can then be estimated using post-processing as described in the proof of Lemma 9. For counting queries and other discrete input we can also use the discrete Gaussian mechanism [CKS20] to estimate $g(X)$, although we might need to change the value of $C$ slightly when $d^{1/4}$ is not an integer. In Section 3.1 we discuss the effect of changing $C$.
@@ 3.1 More accurate estimate of the dataset size
Our mechanism outputs estimates of both $f(X)$ and the size of the dataset $n=|X|$. So far we have ignored the accuracy of the estimate for $n$. However, some tasks require an accurate estimate or $n$. Here we discuss how we can release a more accurate estimate of $n$ and reduce the correlated noise by increasing the magnitude of the independent noise added to each query.
@@
An alternative way of viewing our mechanism is that instead of estimating the value of each query $f(X)_{i}$ independently we privately estimate the difference between $f(X)_{i}$ and $n/2$ for each $i\in[d]$. We can estimate this value more accurately because $f(X)_{i}$ has sensitivity $1$ but $f(X)_{i}-n/2$ only has sensitivity $1/2$. Given an estimate of $n/2$ we can then recover an estimate of $f(X)_{i}$. Let $\tilde{x}$, $\tilde{n}$ denote the output of Algorithm 1. If we subtract $\tilde{n}/2$ from all outputs we find that $\tilde{x}_{i}-\tilde{n}/2=f(X)_{i}-n/2+z_{i}$ is an estimate of the difference between $f(X)_{i}$ and $n/2$ and $\tilde{n}-\tilde{n}/2=n/2+\eta$ is our estimate of $n/2$. The value of $n/2$ itself has a sensitivity of $1/2$. If we computed a new estimate for $n/2$ for each dimension the error would match the standard Gaussian mechanism. The advantage of our approach is that we can reuse the same private estimate of $n/2$ for all $d$ queries.
@@
It is easy to see that the error for the $i$’th query using this representation depends on our estimates of both $n/2$ and $f(X)_{i}-n/2$. We can balance the amount of noise for each estimate which is reflected in the parameter $C$ in the representation of $\hat{X}$ from earlier in this section. The magnitude of noise we use to estimate the value of $f(X)_{i}-n/2$ is $C$ times larger than the noise for estimating $n/2$.
From Lemma 8 we know that $\mathcal{M}(X)\coloneqq g(X)+Z$ satisfies $\mu$-GDP where $g(X)=\sum_{i=1}^{n}\hat{x}_{i}$ and $Z\sim\mathcal{N}(0,\frac{d+C^{2}}{\mu^{2}}I_{d+1})$. Let $B=(d+C^{2})/\mu^{2}$ and $A=B/C^{2}=(d/C^{2}+1^{2})/\mu^{2}$. Then the variance of our estimate of $n/2$, $\mathcal{M}(X)_{d+1}/(2C)$, is $A/4$. If we estimate $f(X)_{i}$ as $\mathcal{M}(X)_{i}/2+\mathcal{M}(X)_{d+1}/(2C)$ the variance of this estimate is $(A+B)/4$. Notice that the dependence of $C$ for the value of $(A+B)/4$ is $(C^{2}+d/C^{2})/4$. Since $C>0$ the value of $(C^{2}+d/C^{2})/4$ is minimized at $C=d^{1/4}$ which is the value we use in Lemma 9 and for Algorithm 1.
@@
Figure 2 shows the effect of varying $C$ when $d=10000$. As stated above the total variance of each query is minimized for $C=d^{1/4}=10$. However, the accuracy of our estimate of $n$ improves as $C$ increases. The plot shows that we can get a much more accurate estimate of $n$ while still outperforming the standard Gaussian mechanism in terms of the total variance. Even if we set $C=d^{1/2}=100$ such that we get an estimate of $n$ with variance only $A=2/\mu^{2}$ the variance for each query is $(d+1)/(2\mu^{2})=5000.5/\mu^{2}$ which is roughly half when compared with $d/\mu^{2}=10000/\mu^{2}$ for the standard Gaussian mechanism. The optimal choice of $C$ depends on the use case.
@@
Figure 2: The plot shows the effect of changing the parameter $C$ for datasets with $d=10000$. On the right, we show the covariance matrix similar to Corollary 11. All entries in the matrix should be scaled by $1/\mu^{2}$. For simplicity we consider $\mu=1$. We have that $A=d/C^{2}+1=10^{4}/C^{2}+1$ and $B=d+C^{2}=10^{4}+C^{2}$. The plot depicts the value of $(A+B)/4$ as a function of $C$.
In this subsection we discussed the trade-off between the private estimate for the size of the input dataset and the error of each query. Throughout this paper we have assumed that we are not given an estimate of $n$ and we explicitly use part of our privacy budget to find such an estimate. However, sometimes we can assume that we have access to an estimate. This is for example the case if the mechanism is used as a subroutine in a bigger system where we already estimated the dataset size. In such cases, we are not required to spend part of our privacy budget on finding a new estimate.
Lemma 12.
@@
It follows from Lemma 4 because the $\ell_{2}$ sensitivity of $g(X)\coloneqq f(X)-n{\bf 1}^{d}/2$ is $\sqrt{d}/2$. ∎
@@ 4 Other settings
In the previous section we presented a simple technique for added elliptical noise for the fundamental task of privately answering $d$ queries from $[0,1]$ under the add/remove neighboring relation. This problem setting was the primary motivation for our work. However, we believe that our technique has the potential to improve the accuracy of differential private mechanisms for other settings as well. To justify this claim we consider an extension of the setting from the previous section. We improve the variance compared to the standard Gaussian mechanism for this problem under both the add/remove and the replacement neighboring relation. We end the section by comparing our technique with a recent result from related work. In Section 6 we discuss potential future extensions.
We consider a setting where a dataset $X$ consists of $n$ data points where $x_{i}\in[0,1]^{m\times d}$. The input is restricted such that each data point has non-zero entries in only a single row. As such we have that $\|x_{i}\|_{2}=(\sum_{j=1}^{m}\sum_{k=1}^{d}x_{i,j,k}^{2})^{1/2}\leq\sqrt{d}$. This kind of simple hierarchical data is is relevant if we are answering $d$ queries and the outputs are grouped based on some property of the data points. Two such examples would be to group users based on location or time when their data was reported. We first consider how the standard Gaussian mechanism estimates $f(X)\coloneqq\sum_{i=1}^{n}x_{i}$ under either neighboring relation.
Lemma 13.
@@
Since $\|x_{i}\|_{2}\leq\sqrt{d}$ the $\ell_{2}$ sensitivity of $f(X)$ is clearly $\sqrt{d}$ under the add/remove neighboring relation. Under the replacement neighboring relation $f(X)$ and $f(X^{\prime})$ differs in at most $2d$ entries by at most $1$ and as such $\|f(X)-f(X^{\prime})\|_{2}\leq\sqrt{2d}$. The privacy guarantees then follow from Lemma 4. ∎
@@
Input : Dataset $X=(x_{1},x_{2},\dots,x_{n})$ where $x_{i}\in[0,1]^{m\times d}$.
Output : $\mu$-GDP estimates of $f(X)\coloneqq\sum_{i=1}^{n}x_{i}$.
@@
4       Sample $\eta_{j}\sim\begin{cases}\mathcal{N}(0,\frac{\sqrt{d}+1}{4\mu^{2}})&\text{under the add/remove neighboring relation,}\\ \mathcal{N}(0,\frac{4}{\mu^{2}})&\text{under the replacement neighboring relation.}\end{cases}$.
@@
6             Sample $z_{j,k}\sim\begin{cases}\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}})&\text{underthe add/remove neighboring relation,}\\ \mathcal{N}(0,\frac{d}{\mu^{2}})&\text{under the replacement neighboring relation.}\end{cases}$.
@@
Algorithm 2 Correlated Gaussian Mechanism when queries are disjoint
Next, we consider how to adapt our mechanism to this setting. The main idea of Algorithm 1 is to add the same noise sample to all queries. This allows us to reduce the magnitude of the independent noise samples for each query when the sensitivity is concentrated along ${\bf 1}^{d}$ as shown for $d=2$ in Figure 1. That is not the case in this setting. Instead we sample correlated noise for each row. We also release an estimate of the number of data points with non-zero entries for each row instead of releasing an estimate of $n$. The pseudo-code is shown in Algorithm 2.
Lemma 14.
@@
In the case of the add/remove relation $f(X)$ and $f(X^{\prime})$ only differ in one row. Algorithm 2 essentially runs Algorithm 1 independently for each row. The privacy guarantees follow from the same proof as Lemma 9.
We prove that the privacy guarantees are correct under replacement using an alternative representation of $X$ similar to Lemma 8 for the setting of Section 3. Define $\hat{x}_{i}\in\mathbb{R}^{m\times(d+1)}$ such that $\hat{x}_{i,j,k}=0$ if $\|x_{i,j}\|_{2}=0$. By definition of the problem setup we have $\|x_{i,j}\|_{2}=0$ for all but one row for each $i\in[n]$. For the row $j\in[m]$ where $\|x_{i,j}\|_{2}>0$ we define $\hat{x}_{i,j,k}=2\cdot x_{i,j,k}-1$, and $\hat{x}_{i,j,d+1}=d^{1/2}$. We define the function $g(X)\coloneqq\sum_{i=1}^{n}\hat{x}_{i}$.
Let $x_{i}$ denote the data point we replace with $x^{\prime}_{i}$ to obtains $X^{\prime}$ and fix the row $j\in[m]$ such that $\|x_{i,j}\|_{2}>0$. We have two cases to consider. (1) If the non-zero entries of $x^{\prime}_{i}$ are also in row $j$ (that is, $\|x^{\prime}_{i,j}\|_{2}>0$) then $\hat{x}_{i,j,k}$ and $\hat{x}^{\prime}_{i,j,k}$ differ by at most $2$ for each $k\in[d]$ and all other entries of $\hat{x}_{i}$ are unchanged. As such $\|g(X)-g(X^{\prime})\|_{2}\leq\sqrt{d\cdot 2^{2}}=2\sqrt{d}$. (2) If the non-zero entries of $x^{\prime}_{i}$ are in another row $j^{\prime}$ the entries of $g(X)$ and $g(X^{\prime})$ differ only in these two rows. In this case we have that $\|g(X)-g(X^{\prime})\|_{2}\leq(2\cdot(\sum_{k=1}^{d}1^{2}+(d^{1/2})^{2}))^{1/2}=\sqrt{4d}=2\sqrt{d}$.
As such, the $\ell_{2}$ sensitivity of $g$ under the replacement neighboring relation is $2\sqrt{d}$ and we can release $g(X)$ under $\mu$-GDP by adding noise from $\mathcal{N}(0,4d/\mu^{2})$ independently to each entry. The proof that Algorithm 2 is equivalent to post-processing the above mechanism follows the steps of the proof for Lemma 9 applied to each row for $C=d^{1/2}$. ∎

### v3

@@ 3 Algorithm
    By definition of the function $g$ we have that $g(X)_{i}=2\cdot f(X)_{i}-n$ for all $i\in[d]$ and $g(X)_{d+1}=Cn=d^{1/4}n$. Let $\tilde{x},\tilde{n}$ denote the output of $\mathcal{M}(X)$. Then for $i\in[d]$ we estimate the $i$’th value of $g(X)$ as
    $h(\mathcal{M}(X))_{i}=2\cdot\tilde{x}_{i}-\tilde{n}=2\cdot(f(X)_{i}+\eta+z_{i})-(n+2\eta)=g(X)+2z_{i}\,.$
    Since $z_{i}\sim\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}})$ we have that $2\cdot z_{i}$ and $Z_{i}$ are both distributed as $\mathcal{N}(0,\frac{d+C^{2}}{\mu^{2}})=\mathcal{N}(0,\frac{d+\sqrt{d}}{\mu^{2}})$. We can estimate the final entry of $g(X)$ as
    $h(\mathcal{M}(X))_{d+1}=d^{1/4}\tilde{n}=d^{1/4}(n+2\eta)=g(X)_{d+1}+d^{1/4}2\eta\,.$
    Both $d^{1/4}2\eta$ and $Z_{d+1}$ are distributed as $\mathcal{N}(0,\frac{d+\sqrt{d}}{\mu^{2}})$. Since the noise terms of all these estimates are independent $h(\mathcal{M}(X))$ is distributed as $\mathcal{M}^{\prime}(X)$.
    Similarly, if $\tilde{y}$ denotes the output of $\mathcal{M}^{\prime}(X)$ we can estimate $f(X)$ and $n$ using the inverse of $h$. We estimate the value of $n$ as
    $h^{-1}(\mathcal{M}^{\prime}(X))_{d+1}=\tilde{y}_{d+1}/d^{1/4}=(g(X)_{d+1}+Z_{d+1})/d^{1/4}=n+Z_{d+1}/d^{1/4}\,,$
    and we estimate the $i$’th entry of $f(X)$ as
    $h^{-1}(\mathcal{M}^{\prime}(X))_{i}=\tilde{y}_{i}/2+\tilde{y}_{d+1}/(2d^{1/4})=(g(X)_{i}+Z_{i})/2+(g(X)_{d+1}+Z_{d+1})/(2d^{1/4})=f(X)_{i}+Z_{i}/2+Z_{d+1}/(2d^{1/4})\,,$
    where $Z_{i}/2$ is distributed as $\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}})$. As such, $h^{-1}(\mathcal{M}^{\prime}(X))$ has the same distribution as $\mathcal{M}(X)$ where $Z_{d+1}/(2d^{1/4})=\eta$ is the correlated noise added to all entries in Algorithm 1.
@@
    The mechanism is Algorithm 1. The error guarantees follow from Lemma 3.1. The privacy guarantees follow from Lemmas 3.3, and 3.4.
Algorithm 1 relies on $d+1$ independent samples from Gaussian distributions. Alternatively, we can view the mechanism as outputting an estimate of $f(X)$ and $n$ by sampling a $d+1$ dimensional multivariate Gaussian distribution. The entries of the covariance matrix contain one of $4$ distinct values depending on (1) whether or not it is on the diagonal and (2) whether or not it is in the final row or column. This representation might be useful for data analysts for estimating the effect of noise. We use this representation again in Section 3.1 where we discuss different values for the parameter $C$.
Corollary 3.1
Let $X$, $n$, and $f(X)$ be defined as in Theorem 3.1 except that $f(X)_{d+1}=n$. Then the mechanism that outputs $f(X)+Z$ where $Z\sim\mathcal{N}(0,\Sigma/\mu^{2})$ satisfies $\mu$-GDP under the add/remove neighboring relation where the covariance matrix $\Sigma\in\mathbb{R}^{(d+1)\times(d+1)}$ has the following structure
@@
In this section we have presented multiple representations of our mechanism. In Lemma 3.4 we showed that Algorithm 1 is equivalent to applying the standard Gaussian mechanism to a different representation of the dataset which we denote $\hat{X}$. However, both of these mechanisms are defined for arbitrary real-valued input and output, and as such any implementation of either representation must be an approximation of the idealized algorithm.
Real values are often approximated in computers using floating-point representation and it is well-known that naive implementations can break the formal privacy guarantees [Mir12]. Additive noise mechanisms such as the Laplace mechanism [DMNS06] and the Gaussian mechanism [DKM+06] are typically defined as releasing the sum of a non-private query output and a sample from a distribution. This can be problematic if the mechanism is implemented using floating-point arithmetics because the output domains might differ between neighboring datasets (see [HDH+22, Theorem 2]). To mitigate this issue [HDH+22] present a technique for mechanisms that add a single noisy value to the true output.
However, in Line 1 of Algorithm 1 we add two noise samples to the query output which likely causes similar issues if floating-point representations are used. To the best of our knowledge, it has not been studied whether implementations of mechanisms that rely on elliptical Gaussian noise suffer from floating-point vulnerabilities. As such, implementing the version of our mechanism from Lemma 3.3 might be preferred for production code. The values of $f(X)$ and $n$ can then be estimated using post-processing as described in the proof of Lemma 3.4. For counting queries and other discrete input we can also use the discrete Gaussian mechanism [CKS20] to estimate $g(X)$, although we might need to change the value of $C$ slightly when $d^{1/4}$ is not an integer or rational number. In Section 3.1 we discuss the effect of the value of $C$ on the private estimates.
@@ 3.1 More accurate estimate of the dataset size
Our mechanism outputs estimates of both $f(X)$ and the size of the dataset $n=|X|$. So far we have ignored the accuracy of the estimate for $n$. However, some tasks require an accurate estimate or $n$. Here we discuss how we can release a more accurate estimate of $n$ and reduce the correlated noise at the cost of increasing the magnitude of the independent noise added to each query.
@@
An alternative way of viewing our mechanism is that instead of estimating the value of each query $f(X)_{i}$ independently, we privately estimate the difference between $f(X)_{i}$ and $n/2$ for each $i\in[d]$. We can estimate this value more accurately because $f(X)_{i}$ has sensitivity $1$ but $f(X)_{i}-n/2$ only has sensitivity $1/2$. Given an estimate of $n/2$, we can then recover an estimate of $f(X)_{i}$. Let $\tilde{x}$, $\tilde{n}$ denote the output of Algorithm 1. If we subtract $\tilde{n}/2$ from all outputs we find that $\tilde{x}_{i}-\tilde{n}/2=f(X)_{i}-n/2+z_{i}$ is an estimate of the difference between $f(X)_{i}$ and $n/2$ and $\tilde{n}-\tilde{n}/2=n/2+\eta$ is our estimate of $n/2$. The value of $n/2$ itself has a sensitivity of $1/2$. If we computed a new estimate for $n/2$ for each query the error would match the standard Gaussian mechanism. The advantage of our approach is that we can reuse the same private estimate of $n/2$ for all $d$ queries.
@@
It is easy to see that the error for the $i$’th query using this representation depends on our estimates of both $f(X)_{i}-n/2$ and $n/2$. We can balance the amount of noise for each estimate which is reflected in the parameter $C$ in the representation of $\hat{X}$ introduced earlier in this section. The magnitude of noise we use to estimate the value of $f(X)_{i}-n/2$ is exactly $C$ times larger than the noise for estimating $n/2$. The following lemma gives the error for the final estimate of $f(X)_{i}$.
Lemma 3.5
Let $\mathcal{M}(X)\coloneqq g(X)+Z$ where $g(X)=\sum_{i=1}^{n}\hat{x}_{i}$ and $Z\sim\mathcal{N}(0,\frac{d+C^{2}}{\mu^{2}}I_{d+1})$. Then $\mathcal{M}$ satisfies $\mu$-GDP under the add/remove neighboring relation. Furthermore, if we estimate the entries of $f(X)$ by post-processing the output of $\mathcal{M}(X)$ similar to Lemma 3.4 the total variance of the noise for each query is $(d+C^{2}+d/C^{2}+1)/(4\mu^{2})$.
  • Proof.
    The privacy guarantee follows from Lemma 3.3. Let $B=(d+C^{2})/\mu^{2}$ and $A=B/C^{2}=(d/C^{2}+1)/\mu^{2}$. Then the variance of our estimate of $n/2$ (that is $\mathcal{M}(X)_{d+1}/(2C)$) is $A/4$. The variance of $f(X)_{i}-n/2$ (that is $\mathcal{M}(X)_{i}/2$) is $B/4$. If we estimate $f(X)_{i}$ as $\mathcal{M}(X)_{i}/2+\mathcal{M}(X)_{d+1}/(2C)$ then the variance is $(A+B)/4$.
Notice that the dependence on $C$ for the variance in Lemma 3.5 is $(C^{2}+d/C^{2})/4$. Since $C>0$, the global minimum of $(C^{2}+d/C^{2})/4$ is at $C=d^{1/4}$ which is the value we use in Lemma 3.4 and for Algorithm 1. The variable $A$ used in the proof is decreasing in $C$ which implies that we can get a more accurate estimate of the dataset size by increasing the value of $C$. However, this will in turn increase the error of the estimate of $f(X)_{i}-n/2$ because the variable $B$ is increasing in $C$.
@@
Figure 2 shows the effect of varying $C$ when $d=10000$. As stated above the total variance of each query is minimized for $C=d^{1/4}=10$. However, the accuracy of our estimate of $n$ improves as $C$ increases. The plot shows that we can get a much more accurate estimate of $n$ while still outperforming the standard Gaussian mechanism in terms of the total variance. Even if we set $C=d^{1/2}=100$ such that we get an estimate of $n$ with variance only $A=2/\mu^{2}$ the variance for each query is $(d+1)/(2\mu^{2})=5000.5/\mu^{2}$ which is roughly half when compared with $d/\mu^{2}=10000/\mu^{2}$ for the standard Gaussian mechanism.
For many problems the choice of $C=d^{1/4}$ is preferred because it gives the best estimate of $f(X)_{i}$. We get no benefit from setting the parameter $C<d^{1/4}$ because the increased magnitude of the correlated noise results in worse estimates both of $n$ and $f(X)_{i}$. However, if we want a better estimate of the dataset size and reduce the magnitude of correlated noise we can increase the value of $C$. This can be beneficial for tasks that heavily rely an the accuracy of the estimate for $n$. It can also be useful if the downstream impact of correlated noise is either difficult to analyze or control. Due to diminishing returns we likely never want to use a value of $C$ larger than $\sqrt{d}$. By setting $C=\sqrt{d}$ we already get an estimate of the dataset size that is independent of $d$ as discussed above. The optimal choice of $C$ depends on the exact use case.
@@
Figure 2: The plot shows the effect of changing the parameter $C$ for datasets with $d=10000$. On the right, we show the covariance matrix similar to Corollary 3.1. All entries in the matrix should be scaled by $1/\mu^{2}$. For simplicity we consider $\mu=1$. We have that $A=d/C^{2}+1=10^{4}/C^{2}+1$ and $B=d+C^{2}=10^{4}+C^{2}$. The plot depicts the value of $(A+B)/4$ as a function of $C$.
In this subsection we discussed the trade-off between the private estimate for the size of the input dataset and the error of each query. Throughout this paper we have assumed that we are not given an estimate of $n$ and we explicitly use part of our privacy budget to find such an estimate. However, sometimes we actually have access to an estimate of $n$. This is for example the case if the mechanism is used as a subroutine in a bigger system where we already computed a private estimated the dataset size. In such cases, we are not required to spend any privacy budget on finding a new estimate.
Lemma 3.6
@@
    It follows from Lemma 2.1 because the $\ell_{2}$ sensitivity of $g(X)\coloneqq f(X)-n{\bf 1}^{d}/2$ is $\sqrt{d}/2$.
@@ 4 Other settings
In the previous section we presented a simple technique for added elliptical noise for the fundamental task of privately answering $d$ queries from $[0,1]$ under the add/remove neighboring relation. This problem setting was the primary motivation for our work. However, we believe that our technique has the potential to improve the accuracy of differentially private mechanisms for other settings as well. To justify this claim we consider an extension of the setting from the previous section. We improve the variance compared to the standard Gaussian mechanism for this problem under both the add/remove and the replacement neighboring relation. We end the section by comparing our technique with a recent result from related work. In Section 6 we discuss potential future extensions.
We consider a setting where a dataset $X$ consists of $n$ data points where $x_{i}\in[0,1]^{m\times d}$. The input is restricted such that each data point has non-zero entries in only a single row, that is |$|\{j:x_{i,j}\neq{\textbf{0}}^{d}\}|\leq 1$. As such we have that $\|x_{i}\|_{2}=(\sum_{j=1}^{m}\sum_{k=1}^{d}x_{i,j,k}^{2})^{1/2}\leq\sqrt{d}$. This kind of simple hierarchical data is relevant if we are answering $d$ queries and the outputs are grouped based on some property of the data points. Two such examples would be to group users based on location or time when their data was reported. We first consider how the standard Gaussian mechanism estimates $f(X)\coloneqq\sum_{i=1}^{n}x_{i}$ under either neighboring relation.
Lemma 4.1
@@
    Since $\|x_{i}\|_{2}\leq\sqrt{d}$ the $\ell_{2}$ sensitivity of $f(X)$ is clearly $\sqrt{d}$ under the add/remove neighboring relation. Under the replacement neighboring relation $f(X)$ and $f(X^{\prime})$ differs in at most $2d$ entries by at most $1$ and as such $\|f(X)-f(X^{\prime})\|_{2}\leq\sqrt{2d}$. The privacy guarantees then follow from Lemma 2.1.
@@
Input : Dataset $X=(x_{1},x_{2},\dots,x_{n})$ where $x_{i}\in[0,1]^{m\times d}$ and for each $i\in[n]$ it holds that |$|\{j:x_{i,j}\neq{\textbf{0}}^{d}\}|\leq 1$.
Output : $\mu$-GDP estimates of $f(X)\coloneqq\sum_{i=1}^{n}x_{i}$ and the count for each row.
@@
4       Sample $\eta_{j}\sim\begin{cases}\mathcal{N}(0,\frac{\sqrt{d}+1}{4\mu^{2}})&\text{under the add/remove neighboring relation,}\\ \mathcal{N}(0,\frac{4}{\mu^{2}})&\text{under the replacement neighboring relation.}\end{cases}$
@@
6             Sample $z_{j,k}\sim\begin{cases}\mathcal{N}(0,\frac{d+\sqrt{d}}{4\mu^{2}})&\text{underthe add/remove neighboring relation,}\\ \mathcal{N}(0,\frac{d}{\mu^{2}})&\text{under the replacement neighboring relation.}\end{cases}$
@@
Algorithm 2 Correlated Gaussian Mechanism for disjoint queries
Next, we consider how to adapt our mechanism to this setting. The main idea of Algorithm 1 is to add the same noise sample to all queries. This allows us to reduce the magnitude of the independent noise samples for each query when the sensitivity is concentrated along ${\bf 1}^{d}$ as shown for $d=2$ in Figure 1. In this setting we do not benefit much from adding the same noise sample to all queries. Instead we sample correlated noise for each row. We also release an estimate of the number of data points with non-zero entries for each row instead of releasing an estimate of $n$. The pseudo-code is shown in Algorithm 2. We prove the privacy guarantees next.
Lemma 4.2
@@
    In the case of the add/remove relation $f(X)$ and $f(X^{\prime})$ only differ in one row. Algorithm 2 essentially runs Algorithm 1 independently for each row. The privacy guarantees follow from the same proof as Lemma 3.4.
    We prove that the privacy guarantees holds under replacement using an alternative representation of $X$ similar to Lemma 3.3 for the setting of Section 3. Define $\hat{x}_{i}\in\mathbb{R}^{m\times(d+1)}$ such that $\hat{x}_{i,j,k}=0$ for all $k\in[d+1]$ if $\|x_{i,j}\|_{2}=0$. By definition of the problem setup we have $\|x_{i,j}\|_{2}=0$ for all but one row for each $i\in[n]$. For the row $j\in[m]$ where $\|x_{i,j}\|_{2}>0$ we define $\hat{x}_{i,j,k}=2\cdot x_{i,j,k}-1$, and $\hat{x}_{i,j,d+1}=d^{1/2}$. We define the function $g(X)\coloneqq\sum_{i=1}^{n}\hat{x}_{i}$.
    Let $x_{i}$ denote the data point we replace with $x^{\prime}_{i}$ to obtain $X^{\prime}$ from $X$ and fix the row $j\in[m]$ such that $\|x_{i,j}\|_{2}>0$. We have two cases to consider (1) If the non-zero entries of $x^{\prime}_{i}$ are also in row $j$ (that is, $\|x^{\prime}_{i,j}\|_{2}>0$) then $\hat{x}_{i,j,k}$ and $\hat{x}^{\prime}_{i,j,k}$ differ by at most $2$ for each $k\in[d]$ for row $j$ and all other entries of $\hat{x}_{i}$ are unchanged. As such we have $\|g(X)-g(X^{\prime})\|_{2}=\|\hat{x}_{i,j}-\hat{x}^{\prime}_{i,j}\|_{2}\leq\sqrt{d\cdot 2^{2}}=2\sqrt{d}$. (2) If the non-zero entries of $x^{\prime}_{i}$ are in another row $j^{\prime}$ the entries of $g(X)$ and $g(X^{\prime})$ differ only in these two rows. Each entry is zero for at least one of $\hat{x}_{i}$ and $\hat{x}^{\prime}_{i}$. In this case we have that $\|g(X)-g(X^{\prime})\|_{2}=\|\hat{x}_{i}-\hat{x}^{\prime}_{i}\|_{2}\leq(2\cdot(\sum_{k=1}^{d}1^{2}+(d^{1/2})^{2}))^{1/2}=\sqrt{4d}=2\sqrt{d}$.
    As such, the $\ell_{2}$ sensitivity of $g$ under the replacement neighboring relation is $2\sqrt{d}$ and we can release $g(X)$ under $\mu$-GDP by adding noise from $\mathcal{N}(0,4d/\mu^{2})$ independently to each entry. The proof that Algorithm 2 is equivalent to post-processing the above mechanism follows the steps of the proof for Lemma 3.4 applied to each row for $C=d^{1/2}$.